-3t - 5= 6 + 7t
A)4
B)10/11
C)-10/11
D)-11/10

Answer by stanbon(48575)

-3t - 5= 6 + 7t
-10t = 11
t = 11/10
--------------------
Cheers,
Stan H.
=================
A)4
B)10/11
C)-10/11
D)-11/10